OpenStudy (anonymous):

Prove that : 4cos2A*(cosA)^2=8(sinA)^2

OpenStudy (lin.ivory):

i have been doing similar questions for the whole day, try using double-angle identities (cosA)^2 =(cos^2 A) and that would be the same thing as (sinA)^2

OpenStudy (anonymous):

I did that but unfortunately its useless ,i didn't reach the proof :(

myininaya (myininaya):

This is an identity we might use: \[\cos(2A)=\cos^2(A)-\sin^2(A)=(1-\sin^2(A))-\sin^2(A)=1-2\sin^2(A)\] \[4(1-2\sin^2(A))\cos^2(A)=4(1-2\sin^2(A)(1-\sin^2(A))= \] \[4(1-\sin^2(A)-2\sin^2(A)+\sin^4(A)) \] \[4(1-3\sin^2(A)+\sin^4(A))\] What you have isn't an identity

myininaya (myininaya):

This is what I'm reading for what you have \[4\cos(2A)(\cos(A))^2=8(\sin(A))^2\] Is that what it said?

OpenStudy (anonymous):

yes ,thats right @myininaya

myininaya (myininaya):

We know this isn't an identity because if A=0 then we have 4=0 which is not true

myininaya (myininaya):

cos(0)=1 sin(0)=0

myininaya (myininaya):

\[4\cos(2 \cdot 0) (\cos(0))^2=8(\sin(0))^2\] \[4\cos(0)(\cos(0))^2=8(\sin(0))^2\] \[4(1)(1)^2=8(0)^2\] \[4=0 \]

myininaya (myininaya):

not true so not an identity

OpenStudy (lin.ivory):

but how do you know that A=0??

myininaya (myininaya):

For this to be an identity it is suppose to be true for all values of A There was a value of A that doesn't satisfy the equality so therefore it is not an identity This is called a counterexample

OpenStudy (lin.ivory):

OHH!! ok! so for future questions, i can use any number for A to prove and check whether the identity is true or not?

myininaya (myininaya):

you cannot use it to prove but you use it to disprove

OpenStudy (anonymous):

so @myininaya ,whats da answer cuz i got confused .*_*

myininaya (myininaya):

You don't understand that I plugged in 0 for A to show the identity is in fact not an identity

myininaya (myininaya):

?

OpenStudy (anonymous):

OHHH! ,ahaaa i got it

myininaya (myininaya):

There are other values of A you could use to disprove also I just chose A because it was the prettiest

myininaya (myininaya):

But Eyad I still think what you have written is wrong

myininaya (myininaya):

Since it says to prove

myininaya (myininaya):

But I could be wrong about you being wrong

myininaya (myininaya):

lol

OpenStudy (anonymous):

lol,but in case thats its written right ,U have just said that we can't use that law (a=0) to prove ,just use it to disprove right ?

OpenStudy (anonymous):

@myininaya : i got another question looks like this one ,maybe u can get a hint from it ,u want me to post it ?

OpenStudy (anonymous):

no, @myininaya , i used the graphing calculator to check their graphs... it is not an identity.

myininaya (myininaya):

@dpalnc Yes I know it is not an identity. I'm saying what he has made not be written correctly since it does say to prove it is an identity.